# Iterated sampling importance resampling with adaptive number of proposals

### v1

@@ 6.3. Properties for the continuous relaxation
$P_{\lambda}(x,A)=\beta(\lambda)P_{\lfloor\lambda\rfloor}(x,A)+(1-\beta(\lambda))P_{\lfloor\lambda\rfloor+1}(x,A),$
@@
(10) $\begin{split}\langle f|P_{\lambda}f\rangle_{\pi}&=\beta(\lambda)\langle f|P_{\lfloor\lambda\rfloor}f\rangle_{\pi}+(1-\beta(\lambda))\langle f|P_{\lfloor\lambda\rfloor+1}f\rangle_{\pi}\\ \epsilon(\lambda,x)&=\beta(\lambda)\epsilon(\lfloor\lambda\rfloor,x)+(1-\beta(\lambda))\epsilon(\lfloor\lambda\rfloor+1,x)\\ \epsilon(\lambda)&=\beta(\lambda)\epsilon(\lfloor\lambda\rfloor)+(1-\beta(\lambda))\epsilon(\lfloor\lambda\rfloor+1),\end{split}$
@@
Let $f\in L^{2}(\pi)$, then $\lambda\mapsto\langle f|P_{\lambda}f\rangle_{\pi}$, $\lambda\geq 1$ is continuous non-negative decreasing convex function. Additionally we have the following properties:
@@
Follows from Lemma 44 in Appendix D, when combined with Lemma 13, 15 and 16. ∎
@@
By definition $\epsilon(N)$ and $\epsilon(N,x)$ are positive for all $N\in\mathbb{N}_{+}$. Therefore, the claim follows from Lemma 44 in Appendix D, when combined with Lemma 18. ∎
@@
Since $x$ is an atom we have $P_{1}(x,\{x\})=\langle g_{x}|g_{x}\rangle_{\pi}$ and if $N\in\mathbb{N}_{+}$, $N\geq 2$, then
@@
Thus from (6) we get the claim for $N\in\mathbb{N}_{+}$, from which the general version follows. ∎
@@
When $x$ is not an atom the claim follows from Theorem 21. When $x$ is an atom the claim follows from Lemma 22 and Theorem 20 by noticing $\bar{g}_{x}\in L_{0,+}^{2}(\pi)$. ∎
@@
Function $\epsilon:(1,\infty)\to[0,1)$, $\epsilon(\lambda)$ is positive convex and strictly decreasing by Theorem 21. Thus by noticing that $g:[0,1)\to\mathbb{R}$, $g(x)=\frac{1+x}{1-x}$ is positive strictly increasing strictly convex function, we get that $g(\epsilon(\lambda))$ is positive strictly decreasing function and it holds for $1<\lambda_{1}<\lambda_{2}$ and $t\in(0,1)$ that
$g(\epsilon(t\lambda_{1}+(1-t)\lambda_{2}))\leq g(t\epsilon(\lambda_{1})+(1-t)\epsilon(\lambda_{2}))<tg(\epsilon(\lambda_{1}))+(1-t)g(\epsilon(\lambda_{2})).\qed$
@@
Function $\lambda\mapsto\int_{\mathbb{S}_{0}}\epsilon(\lambda,x)\pi(\mathrm{d}x)+\int_{\mathbb{S}_{+}}\pi(\{x\})P_{\lambda}(x,\{x\})\in(\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x),1]$ is positive convex and strictly decreasing by Theorem 21, 23 and Lemma 19. Therefore we obtain the claim for $\psi$ by noticing $g:(\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x),1]\to(0,1]$, $g(y)=\frac{y-\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x)}{1-\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x)}$ is linear function. Since $\psi(1)=1$, we have $\psi(y)\in[0,1)$ for all $y\in(1,\infty)$, hence rest of the proof is similar to the proof of Lemma 24. ∎
@@ 7. Asymptotic variance
when choosing $x=0.2$ and comparing $P_{2}(x,\;\cdot\;)$ and $P_{3}(x,\;\cdot\;)$ with both sets $(0,0.1)$ and $(0.9,1)$. However there is a weaker condition for showing monotonicity of asymptotic variance, namely the covariance ordering [tierney-note, mira-geyer], which in our setting means the monotonicity of $\lambda\mapsto\langle f|P_{\lambda}f\rangle_{\pi}$ for all $f\in L^{2}(\pi)$. In this paper monotonicity and convexity of asymptotic variance are obtained in a similar fashion as in [e.g. andrieu-vihola-order, Proposition 26] using Dirichlet forms, these results are combined in Theorem 59 in Appendix H. Without loss of generality we can study $L_{0}^{2}(\pi)$ functions.
@@
We have $P_{N}$ is $\pi$-reversible for $N\in\mathbb{N}_{+}$, $N\geq 2$ [e.g. tjelmeland, andrieu-lee-vihola], which also holds for $N=1$, since $P_{1}(x,\mathrm{d}y)=\delta_{x}(\mathrm{d}y)$. Therefore we have $P_{\lambda}$ is $\pi$-reversible for $\lambda\geq 1$, since
@@
where $c=\frac{\lambda-1}{2\hat{w}+\lambda-1}$ and $Q$ is $\pi$-reversible Markov kernel. Thus for $f\in L_{0}^{2}(\pi)$
@@
since the Dirichlet form is non-negative for Markov kernel. Therefore $I-P_{\lambda}$ is invertible by Lemma 55 in Appendix H. ∎
@@
The claim follows from Theorem 59 in Appendix H when combined with Lemma 26 and Theorem 20. ∎
@@
Let $f\in L_{0,+}^{2}(\pi)$, then for Algorithm 2, it holds that $\lambda\mapsto\mathcal{E}_{\lambda}(f)$, $\lambda\geq 1$ is non-negative, continuous, concave and strictly increasing function.
@@
Since $\mathcal{E}_{\lambda}(g)=\langle f|f\rangle_{\pi}-\langle f|P_{\lambda}f\rangle_{\pi}$ and $\langle f|P_{1}f\rangle_{\pi}=\langle f|f\rangle$, the claim follows from Theorem 20. ∎
@@ 8.1. Rejection probabilities
By definition both $\epsilon(\lambda,z_{1})$ and $\epsilon(\lambda)$ are non-negative. Let $N\in\{2,3,\dots\}$
@@
where the first inequality follows from Jensen’s inequality. Therefore
$\epsilon(N)=\int_{\mathbb{S}}\epsilon(N,z_{1})\pi(\mathrm{d}z_{1})\leq\frac{2\hat{w}-1}{2\hat{w}+N-2}$
@@
$\epsilon(\lambda,z_{1})\leq\frac{2\hat{w}}{2\hat{w}+\lambda-1},$
$\epsilon(\lambda)\leq\frac{2\hat{w}}{2\hat{w}+\lambda-1}.$
@@
$\epsilon(N)$ $\geq\int_{\mathbb{S}}\frac{w(z_{1})}{w(z_{1})+N-1}\pi(\mathrm{d}z_{1})=\int_{\mathbb{X}}\frac{w(z_{1})^{2}}{w(z_{1})+N-1}q(\mathrm{d}z_{1})$
$=\int_{\mathbb{X}}g(w(z_{1}))q(\mathrm{d}z_{1})\geq\frac{1}{N},$
where the last inequality follows from Jensen’s inequality, since
@@
is convex. Thus for $\lambda\geq 1$ it holds by Lemma 46 in Appendix D
@@
where $\beta(\lambda)=\lfloor\lambda\rfloor+1-\lambda\in[0,1)$. We shall define the function on the right hand side of (11), which was already mentioned in Section 5: For Algorithm 2 the upper bound function of squared rejection probability obtained by Jensen’s inequality is denoted by $\epsilon_{s}:[1,\infty)\to\mathbb{R}$,
@@
It holds that $\epsilon_{s}:[1,\infty)\to\mathbb{R}$ is strictly decreasing positive function and $\epsilon(\lambda)^{2}\leq\epsilon_{s}(\lambda)$. If Assumption (A1) holds $\epsilon_{s}(\lambda)\leq\lfloor\lambda\rfloor^{-1}\hat{w}$.
@@
First inequality follows from (11). When $N\in\mathbb{N}_{+}$ and $N\leq\lambda_{1}<\lambda_{2}<N+1$, then by the definition of $\epsilon_{s}$ we have $\epsilon_{s}(\lambda_{1})>\epsilon_{s}(\lambda_{2})$. If $\lambda\in[N,N+1)$, then
@@
$=\int_{\mathbb{W}_{N+1}}w(z_{1})^{3}\bigg(\frac{\beta(\lambda)w(z_{N+1})+\sum_{k=1}^{N}w(z_{k})}{(\sum_{j=1}^{N}w(z_{j}))(\sum_{j=1}^{N+1}w(z_{j}))}\bigg)^{2}-w(z_{1})^{3}\bigg(\frac{1}{\sum_{i=1}^{N+1}w(z_{i})}\bigg)^{2}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)$
$=\int_{\mathbb{W}_{N+1}}w(z_{1})^{3}\frac{(\beta(\lambda)w(z_{N+1})+\sum_{k=1}^{N}w(z_{k}))^{2}-(\sum_{i=1}^{N}w(z_{j}))^{2}}{(\sum_{i=1}^{N}w(z_{j}))^{2}(\sum_{l=1}^{N+1}w(z_{l}))^{2}}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)$
$=\int_{\mathbb{W}_{N+1}}w(z_{1})^{3}\frac{\beta(\lambda)w(z_{N+1})[\beta(\lambda)w(z_{N+1})+2\sum_{k=1}^{N}w(z_{k})]}{(\sum_{i=1}^{N}w(z_{j}))^{2}(\sum_{l=1}^{N+1}w(z_{l}))^{2}}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)>0.$
Therefore $\epsilon_{s}$ is strictly decreasing. When Assumption (A1) holds we have
@@
Let us define for Algorithm 2 the derivative of average rejection probability. The derivative function $\epsilon^{\prime}$ is undefined for $\lambda\in\mathbb{N}_{+}$, which can be seen from (10) and Lemma 18. We will use the same notation for the extended derivative function
@@
since they are the same everywhere the derivative is defined. The following bounds hold for this function:
@@
The extended derivative function $\epsilon^{\prime}(\lambda):[1,\infty)\to\mathbb{R}$ is negative function and when $N\in\mathbb{N}_{+}$ it holds
$\epsilon^{\prime}(x)$ $=\epsilon^{\prime}(y)\qquad\text{for $N\leq x\leq y<N+1$,}$
$\epsilon^{\prime}(x)$ $<\epsilon^{\prime}(y)\qquad\text{for $N\leq x<N+1\leq y$}.$
If Assumption (A1) holds
@@
where $C=\int_{\mathbb{S}}w(x)\pi(\mathrm{d}x)\in(0,\hat{w}$].
@@
Let us assume first Assumption (A1) holds. Let $N\in\mathbb{N}_{+}$ and let us denote $B_{n}=\sum_{i=1}^{n}w(z_{i})$, then
@@
where the last inequality follows from Jensen’s inequality. Rest of the claims follow from Lemma 18 and the definition of $\epsilon^{\prime}$. ∎
@@
We have $P_{\lambda}(x,\{x\})>\pi(\{x\})$ by Lemma 19 and Theorem 23. Rest of the inequalities follow from each other and Lemma 48 in Appendix E, when $A=\{x\}$. ∎

### v2

@@ 6.3. Properties for the continuous relaxation
$P_{\lambda}(x,A)=\beta(\lambda)P_{\lfloor\lambda\rfloor}(x,A)+\big(1-\beta(\lambda)\big)P_{\lfloor\lambda\rfloor+1}(x,A),$
@@
(10) $\begin{split}\langle f|P_{\lambda}f\rangle_{\pi}&=\beta(\lambda)\langle f|P_{\lfloor\lambda\rfloor}f\rangle_{\pi}+\big(1-\beta(\lambda)\big)\langle f|P_{\lfloor\lambda\rfloor+1}f\rangle_{\pi}\\ \epsilon(\lambda,x)&=\beta(\lambda)\epsilon(\lfloor\lambda\rfloor,x)+\big(1-\beta(\lambda)\big)\epsilon(\lfloor\lambda\rfloor+1,x)\\ \epsilon(\lambda)&=\beta(\lambda)\epsilon(\lfloor\lambda\rfloor)+\big(1-\beta(\lambda)\big)\epsilon(\lfloor\lambda\rfloor+1),\end{split}$
@@
Let $f\in L^{2}(\pi)$, then $\lambda\mapsto\langle f|P_{\lambda}f\rangle_{\pi}$, $\lambda\geq 1$ is continuous non-negative decreasing convex function. Additionally, we have the following properties:
@@
Follows from Lemma 44 in Appendix D when combined with Lemma 13, 15 and 16. ∎
@@
By definition $\epsilon(N)$ and $\epsilon(N,x)$ are positive for all $N\in\mathbb{N}_{+}$. Therefore, the claim follows from Lemma 44 in Appendix D when combined with Lemma 18. ∎
@@
Since $x$ is an atom, we have $P_{1}(x,\{x\})=\langle g_{x}|g_{x}\rangle_{\pi}$ and if $N\in\mathbb{N}_{+}$, $N\geq 2$, then
@@
Thus, by (6) we get the claim for $N\in\mathbb{N}_{+}$, from which the general version follows. ∎
@@
When $x$ is not an atom, the claim follows from Theorem 21. When $x$ is an atom, the claim follows from Lemma 22 and Theorem 20 by noticing $\bar{g}_{x}\in L_{0,+}^{2}(\pi)$. ∎
@@
Function $\epsilon:(1,\infty)\to[0,1)$, $\epsilon(\lambda)$ is positive convex and strictly decreasing by Theorem 21. Thus, by noticing that $g:[0,1)\to\mathbb{R}$, $g(x)=\frac{1+x}{1-x}$ is positive strictly increasing strictly convex function, we get that $g(\epsilon(\lambda))$ is positive strictly decreasing function and it holds for $1<\lambda_{1}<\lambda_{2}$ and $t\in(0,1)$ that
$g\big(\epsilon(t\lambda_{1}+(1-t)\lambda_{2})\big)\leq g\big(t\epsilon(\lambda_{1})+(1-t)\epsilon(\lambda_{2})\big)<tg\big(\epsilon(\lambda_{1})\big)+(1-t)g\big(\epsilon(\lambda_{2})\big).\qed$
@@
Function $\lambda\mapsto\int_{\mathbb{S}_{0}}\epsilon(\lambda,x)\pi(\mathrm{d}x)+\int_{\mathbb{S}_{+}}\pi(\{x\})P_{\lambda}(x,\{x\})\in(\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x),1]$ is positive convex and strictly decreasing by Theorem 21, 23 and Lemma 19. Therefore, we obtain the claim for $\psi$ by noticing $g:(\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x),1]\to(0,1]$, $g(y)=\frac{y-\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x)}{1-\int_{\mathbb{S}_{+}}\pi(\{x\})\pi(\mathrm{d}x)}$ is linear function. Since $\psi(1)=1$, we have $\psi(y)\in[0,1)$ for all $y\in(1,\infty)$, hence rest of the proof is similar to the proof of Lemma 24. ∎
@@ 7. Asymptotic variance
when choosing $x=0.2$ and comparing $P_{2}(x,\;\cdot\;)$ and $P_{3}(x,\;\cdot\;)$ with both sets $(0,0.1)$ and $(0.9,1)$. However, there is a weaker condition for showing monotonicity of asymptotic variance, namely the covariance ordering [tierney-note, mira-geyer], which in our setting means the monotonicity of $\lambda\mapsto\langle f|P_{\lambda}f\rangle_{\pi}$ for all $f\in L^{2}(\pi)$. In this paper monotonicity and convexity of asymptotic variance are obtained in a similar fashion as in [e.g. andrieu-vihola-order, Proposition 26] using Dirichlet forms, these results are combined in Theorem 63 in Appendix H. Without loss of generality we can study $L_{0}^{2}(\pi)$ functions.
@@
We have $P_{N}$ is $\pi$-reversible for $N\in\mathbb{N}_{+}$, $N\geq 2$ [e.g. tjelmeland, andrieu-lee-vihola], which also holds for $N=1$, since $P_{1}(x,\mathrm{d}y)=\delta_{x}(\mathrm{d}y)$. Therefore, we have $P_{\lambda}$ is $\pi$-reversible for $\lambda\geq 1$, since
@@
where $c=\frac{\lambda-1}{2\hat{w}+\lambda-1}$ and $Q$ is $\pi$-reversible Markov kernel. Thus, for $f\in L_{0}^{2}(\pi)$
@@
since the Dirichlet form is non-negative for Markov kernel. Therefore, $I-P_{\lambda}$ is invertible by Lemma 59 in Appendix H. ∎
@@
The claim follows from Theorem 63 (v) in Appendix H when combined with Lemma 26 and Theorem 20. ∎
@@
Let $f\in L_{0,+}^{2}(\pi)$, then for Algorithm 2 it holds that $\lambda\mapsto\mathcal{E}_{\lambda}(f)$, $\lambda\geq 1$ is non-negative, continuous, concave and strictly increasing function.
@@
Since $\mathcal{E}_{\lambda}(g)=\langle f|f\rangle_{\pi}-\langle f|P_{\lambda}f\rangle_{\pi}$ and $\langle f|P_{1}f\rangle_{\pi}=\langle f|f\rangle_{\pi}$, the claim follows from Theorem 20. ∎
@@ 8.1. Rejection probabilities
Let $N\in\{2,3,\dots\}$, then
@@
where the third equality follows from exchangeability of $z_{2},\dots,z_{N}$ under the integral and the first inequality from Jensen’s inequality. Therefore
$\epsilon(N)=\int_{\mathbb{S}}\epsilon(N,z_{1})\pi(\mathrm{d}z_{1})\leq\frac{2\hat{w}-1}{2\hat{w}+N-2},$
@@
$\epsilon(\lambda,z_{1})\leq\frac{2\hat{w}}{2\hat{w}+\lambda-1},\qquad\epsilon(\lambda)\leq\frac{2\hat{w}}{2\hat{w}+\lambda-1}.$
@@
$\epsilon(N)$ $\geq\int_{\mathbb{S}}\frac{w(z_{1})}{w(z_{1})+N-1}\pi(\mathrm{d}z_{1})=\int_{\mathbb{X}}\frac{w(z_{1})^{2}}{w(z_{1})+N-1}q(\mathrm{d}z_{1})\geq\frac{1}{N},$
where the last inequality follows from Jensen’s inequality, since $g:\mathbb{R}\to\mathbb{R}$,
@@
is convex. Thus, for $\lambda\geq 1$ it holds by Lemma 46 in Appendix D
@@
where $\beta(\lambda)=\lfloor\lambda\rfloor+1-\lambda\in(0,1]$. We shall define the function on the right-hand side of (11), which was already mentioned in Section 5: For Algorithm 2 the upper bound function of squared average rejection probability obtained by Jensen’s inequality is denoted by $\epsilon_{s}:[1,\infty)\to\mathbb{R}$,
@@
It holds that $\epsilon_{s}:[1,\infty)\to\mathbb{R}$ is strictly decreasing positive function, $\epsilon(\lambda)^{2}\leq\epsilon_{s}(\lambda)$ and $\epsilon_{s}(1)=1$. If Assumption (A1) holds, then $\epsilon_{s}(\lambda)\leq\lfloor\lambda\rfloor^{-1}\hat{w}$.
@@
First inequality follows from (11) and $\epsilon_{s}(1)=1$ by the definition. When $N\in\mathbb{N}_{+}$ and $N\leq\lambda_{1}<\lambda_{2}<N+1$, then by the definition of $\epsilon_{s}$, we have $\epsilon_{s}(\lambda_{1})>\epsilon_{s}(\lambda_{2})$. If $\lambda\in[N,N+1)$, then
@@
$=\int_{\mathbb{W}_{N+1}}w(z_{1})^{3}\bigg(\frac{\beta(\lambda)w(z_{N+1})+\sum_{k=1}^{N}w(z_{k})}{(\sum_{j=1}^{N}w(z_{j}))(\sum_{l=1}^{N+1}w(z_{l}))}\bigg)^{2}-w(z_{1})^{3}\bigg(\frac{1}{\sum_{l=1}^{N+1}w(z_{l})}\bigg)^{2}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)$
$=\int_{\mathbb{W}_{N+1}}w(z_{1})^{3}\frac{(\beta(\lambda)w(z_{N+1})+\sum_{k=1}^{N}w(z_{k}))^{2}-(\sum_{i=1}^{N}w(z_{i}))^{2}}{(\sum_{j=1}^{N}w(z_{j}))^{2}(\sum_{l=1}^{N+1}w(z_{l}))^{2}}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)$
$=\int_{\mathbb{W}_{N+1}}w(z_{1})^{3}\frac{\beta(\lambda)w(z_{N+1})[\beta(\lambda)w(z_{N+1})+2\sum_{k=1}^{N}w(z_{k})]}{(\sum_{j=1}^{N}w(z_{j}))^{2}(\sum_{l=1}^{N+1}w(z_{l}))^{2}}\bigg(\prod_{n=1}^{N+1}q(\mathrm{d}z_{n})\bigg)>0.$
Therefore, $\epsilon_{s}$ is strictly decreasing. When Assumption (A1) holds, we have
@@
Let us define for Algorithm 2 the derivative of average rejection probability. The derivative function $\epsilon^{\prime}$ is undefined for $\lambda\in\mathbb{N}_{+}$, which can be seen from (10) and Lemma 18. We will use the same notation for the extended derivative function:
@@
which is a càdlàg function. The following bounds hold for this function:
@@
The extended derivative function $\epsilon^{\prime}(\lambda):[1,\infty)\to\mathbb{R}$ is increasing negative function. Let $N\in\mathbb{N}_{+}$, then $\epsilon^{\prime}(x)=\epsilon^{\prime}(y)$ if $N\leq x\leq y<N+1$, and $\epsilon^{\prime}(x)<\epsilon^{\prime}(y)$ if $N\leq x<N+1\leq y$. If Assumption (A1) holds, then
@@
where $C=\int_{\mathbb{S}}w(x)\pi(\mathrm{d}x)\in[1,\hat{w}$].
@@
Let us first assume Assumption (A1) holds. Let $N\in\mathbb{N}_{+}$ and let us denote $B_{n}=\sum_{i=1}^{n}w(z_{i})$, then
@@
where the last inequality follows from Jensen’s inequality. Let us define $\nu=\mu|_{\mathbb{S}}$, then by Hölder’s inequality
$\Big(\int_{\mathbb{S}}\Big|\frac{\pi(x)}{\sqrt{q(x)}}\sqrt{q(x)}\Big|\nu(\mathrm{d}x)\Big)^{2}\leq\int_{\mathbb{S}}\Big|\frac{\pi(x)}{\sqrt{q(x)}}\Big|^{2}\nu(\mathrm{d}x)\int_{\mathbb{S}}\Big|\sqrt{q(y)}\Big|^{2}\nu(\mathrm{d}y),$
hence
$\int_{\mathbb{S}}w(x)\pi(\mathrm{d}x)=\int_{\mathbb{S}}\frac{\pi(x)^{2}}{q(x)}\nu(\mathrm{d}x)\geq\frac{\Big(\int_{\mathbb{S}}\pi(x)\nu(\mathrm{d}x)\Big)^{2}}{\int_{\mathbb{S}}q(x)\nu(\mathrm{d}x)}=\frac{1}{\int_{\mathbb{S}}q(x)\mu(\mathrm{d}x)}\geq 1.$
Rest of the claims follow from Lemma 18 and the definition of $\epsilon^{\prime}$. ∎
@@
We have $P_{\lambda}(x,\{x\})>\pi(\{x\})$ by Lemma 19 and Theorem 23. Rest of the inequalities follow from each other and Lemma 49 in Appendix E when $A=\{x\}$. ∎